# Nonequilibrium thermodynamics of populations of weakly-coupled low-temperature-differential Stirling engines with synchronous and asynchronous transitions

### v1

@@ Abstract
This study developed the theory of nonequilibrium thermodynamics for populations of low-temperature-differential (LTD) Stirling engines weakly-coupled in a general class of networks to clarify the effects of synchronous and asynchronous transitions on the power and thermal efficiency. We first show that synchronous (asynchronous) transitions increase (decrease) the power and thermal efficiency of weakly-coupled LTD Stirling engines based on quasilinear response relations between formally defined thermodynamic fluxes and forces. After that, we construct a conceptual model satisfying the quasilinear response relations to give a physical interpretation of the changes in power and thermal efficiency due to synchronous and asynchronous transitions, and justify the use of this conceptual model. We then show that the conceptual model, rather than the quasilinear response relations, preserves the irreversible nature in the relative motion of the original model and thus shows more accurate results than the analysis using the quasilinear response relations. Finally, we compare the dynamics between the original and the conceptual models for two-engine systems and show that the conceptual model roughly preserves the dynamical characteristics leading up to the synchronous transitions, while some detailed dynamical structures are lost.
@@ I Introduction
The present study extends the analysis performed in [9] to populations of LTD Stirling engines weakly-coupled in a general class of networks, and constructs the theory of nonequilibrium thermodynamics with synchronous and asynchronous transitions. In Section II, we give a mathematical model of weakly-coupled low-temperature differential Stirling engines. In Section III, we develop the theory of nonequilibrium thermodynamics for the coupled engines as the main results. We first perform the analysis using formally defined thermodynamic fluxes and forces and the quasilinear response relations established between them to show that synchronous (asynchronous) transitions increase (decrease) the power and thermal efficiency. After that, we construct a conceptual model that satisfies the quasilinear response relations to give a physical interpretation of the transitions in power and thermal efficiency, and justify the use of this conceptual model. We then show that the analysis using the conceptual model, rather than the quasilinear response relations, preserves the irreversible nature in the relative motion and thus leads to more accurate results than those using the quasilinear response relations. In Section IV, we compare the dynamics between the original and the conceptual models for two-engine systems and show that the conceptual model roughly preserves the dynamical characteristics leading up to the synchronous transitions, while some detailed dynamical structures are lost. Finally, in Section V, we summarize the study and give the future outlook.
@@ II Model
Here, $\theta$ is the phase of the crank connected to the power piston, and $\omega$ is the angular velocity of that crank. A crank is also attached to the displacer, which is advanced by $\pi/2$ in phase compared to the crank attached to the power piston. These cranks are attached to a flywheel with a large moment of inertia to smoothen the rotational speed of the engine. The physical meanings of the parameters included in the above model are as follows: $\sigma$ is a positive constant determined by the surface areas of the large and small cylinders; $V(\theta)=2+\sigma(1-\cos\theta)$ and $T(\theta,\omega)=T_{\rm eff}(\theta)/\left(1+\frac{\sigma\sin\theta\omega}{GV(\theta)}\right)$ represent the volume and temperature of the gas confined to the cylinders, respectively; $T_{\rm eff}(\theta)=1+\frac{\sin\theta}{2}\Delta T$ is the effective temperature of the heat reservoirs that periodically changes depending on the phase; $G$ is the thermal conductance associated with the heat transfer between the gas and the surface of the large cylinder; $P_{\rm air}$ is the atmospheric pressure acting on the power piston; $\Gamma$ is the friction coefficient associated with the flywheel, and $T_{\rm load}>0$ is the load torque acting on the cranks. The minimal model was obtained by assuming that the heat fluxes from the bottom and top surfaces of the large cylinder obey the Fourier law
@@
where $G_{\rm m}(\theta)$ with ${\rm m}={\rm b}$ (or t) represents the effective thermal conductance between the gas and the bottom (or top) heat reservoir. It was also assumed that $G_{\rm m}(\theta)\equiv G\chi_{\rm m}(\theta)$, where $\chi_{\rm m}(\theta)\hskip 2.84526pt(0\leq\chi_{\rm m}(\theta)\leq 1)$ is a function that controls the coupling between the gas and the bottom or top heat reservoir, given as $\chi_{\rm b}(\theta)=\frac{1}{2}(1+\sin\theta)$ and $\chi_{\rm t}(\theta)=\frac{1}{2}(1-\sin\theta)$, reflecting the motion of the displacer [5]. The dynamical equations describe the engine as a nonlinear pendulum, where the first term on the right-hand side of Eq. (1b) represents the driven force due to the temperature difference. It has been experimentally verified that the minimal model (1a) and (1b) correctly reflects the essential characteristics of a real LTD Stirling engine [12].
Figure 1: (a) Front view of an LTD Stirling engine. The gas confined to the cylinders is in contact with the bottom and top heat reservoirs with temperature $T_{\rm b}$ and $T_{\rm t}$. The displacer that advances the power piston in $\pi/2$ phase controls the coupling between the gas and the heat reservoirs. (b) Side view of $N$ LTD Stirling engines coupled in a chain with different load torques acting on the cranks.
@@
Some important assumptions are made about the equations of motion (3a) and (3b) describing the rotational motion and the equation (4) describing the heat flux. We first assume that each engine is in the quasilinear response regime [5] in the absence of coupling, i.e., a limit cycle circling the surface of the phase space $\mathbb{T}\times\mathbb{R}$ that describes the rotational motion exists, and the effective frequency $\langle\omega_{i}\rangle$ and the averaged heat flux $\Bigl{\langle}J_{Q_{\rm b}}^{(i)}\Bigr{\rangle}$ from the heat reservoir of temperature $T_{\rm b}$ show a linear dependency w.r.t. $T_{\rm load}^{(i)}$ and $\Delta T$:
@@
Figure 2: Typical trajectories (red and green plots) of the limit cycle motion of the two engines satisfying condition (9). The parameters of the engines were chosen as follows: $\sigma=0.02$, $p_{\mathrm{air}}=\frac{1}{V(\frac{\pi}{4})}=\frac{1}{2+\sigma(1-\cos\theta)}\approx 0.49854$, $G=1.5$, $\Gamma=0.001$, $\Delta T=1/29.3$, $T_{\mathrm{load}}^{(1)}=8.5324\times 10^{-7}$, and $T_{\mathrm{load}}^{(2)}=1.2799\times 10^{-5}$.
@@ III Main results: Theory of power and thermal efficiency based on nonequilibrium thermodynamics
In this section, we develop a theory of nonequilibrium thermodynamics for weakly-coupled LTD Stirling engines with synchronous and asynchronous transitions. We first use the quasilinear response relations between thermodynamic fluxes and forces to show that the synchronous (asynchronous) transitions increase (decrease) the power and thermal efficiency in subsection A. After that, we construct a conceptual model satisfying the above quasilinear response relations to give a physical interpretation of the changes in power and thermal efficiency due to the weak coupling in subsection B. We then discuss the justification for using this conceptual model in subsection C. In subsection D, we perform numerical experiments to confirm the validity of the analysis on power and thermal efficiency. Finally, in subsection E, we review the analysis in subsections A -- C and show that the analysis using the conceptual model, rather than the quasilinear response relations, preserves the irreversible nature in the relative motion and thus leads to more accurate results than those using the quasilinear response relations.
@@ III.1 Analysis based on the quasilinear response relations
Here, $\bigl{\langle}P_{\rm m}\bigr{\rangle}\equiv\bigl{\langle}\omega_{\rm m}\bigr{\rangle}\left(\sum_{i=1}^{N}T_{\rm load}^{(i)}\right)$ denotes the power owing to the motion of the mean angle, and $\bigl{\langle}P_{\rm rel}\bigr{\rangle}\equiv\sum_{i=1}^{N}\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}T_{\rm load}^{(i)}$ denotes the power owing to the relative motion. To see the effects of synchronous and asynchronous transitions on the brake power, it is sufficient to investigate the effects on the power owing to the relative motion $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ since $\bigl{\langle}\omega_{\rm m}\bigr{\rangle}$ is independent of the coupling from the quasilinear response relations. Substituting the quasilinear response relations w.r.t. fluxes $\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$ in Eq. (17), $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ can be rewritten as
@@
We have added a term representing the power due to the weak coupling in Eq. (26), which takes the value 00 and is approximated as
@@
and used the quasilinear response relations w.r.t. fluxes $\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$ in Eq. (27). The first equality in equation (28) can be easily derived by noting that $K_{ij}=K_{ji}$; the second approximation comes from Eq. (14) and the third equality can be obtained using Eq. (16). Since $\frac{1}{N}\sum_{i=1}^{N}\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}^{2}$ represents the dispersion of $\langle\omega_{i}\rangle$ around the mean value $\langle\omega_{\rm m}\rangle$, Eq. (27) implies that the coupled system achieves higher averaged brake power with smaller rotational speed dispersion, and that the averaged brake power reaches the maximum when the engines are synchronized. On the other hand, the total heat flux from the heat reservoir of temperature $T_{\rm b}$ is independent of the coupling from the quasilinear response relations. This implies that synchronous (asynchronous) transitions improve (deteriorate) both the averaged brake power and the thermal efficiency of the total system.
@@ III.2 Analysis based on the conceptual models constructed from the quasilinear response relations
Equations (31)-(34) indicate that system M can be considered as composed of the same heat reservoirs as those of the original system and a heat engine ${\rm M}_{\rm EG}$ working between the heat reservoirs with viscous damping coefficient of $N\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)$ and load torque $\sum_{i=1}^{N}T_{\rm load}^{(i)}$. Indeed, in this case, if we assume that the thermodynamic irreversibility in system M is only due to the heat exchange between the two heat reservoirs and the heat engine ${\mathrm{M}}_{\mathrm{EG}}$, the entropy production rate $\frac{d\sigma_{\mathrm{M}}}{dt}$ is obtained as
$\displaystyle\frac{d\sigma_{\mathrm{M}}}{dt}=-\frac{J_{Q_{\mathrm{b}}}^{\mathrm{M}}}{T_{\mathrm{b}}}-\frac{J_{Q_{\mathrm{t}}}^{\mathrm{M}}-P_{\mathrm{fric}}^{\mathrm{M}}}{T_{\mathrm{t}}}.$ (35)
By utilizing the energy conservation law
@@
$\biggl{\langle}\frac{d\sigma_{\mathrm{D}}}{dt}\biggr{\rangle}$ $\approx\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)\Biggl{\langle}\sum_{i=1}^{N}\left(\omega_{\mathrm{d}}^{(i)}\right)^{2}\Biggr{\rangle}$ (44)
@@
Now we use the conceptual model constructed above to illustrate how weak coupling increases the averaged brake power and thermal efficiency of the total system (Note that the conceptual model is refered to as the physical picture of isolated systems M and D, not only the differential equations describing the rotational motion and the heat flux equation. This model was constructed through the quasilinear response relations (17) and thus should reflects the irreversibility of the original coupled Stirling engine system). Figure 3 shows a schematic diagram of systems M and D. The load acting on the engine ${\rm M}_{\rm EG}$ and flywheels ${\rm D}_{\rm FW}^{(i)}$ are represented by weights tied to the rotating bodies. Let us first consider the case where the original Stirling engines are uncoupled. In this case, flywheels ${\rm D}_{\rm FW}^{(i)}$ do not interact either and are simply driven by load torques of magnitude $\left|T_{\rm load}^{(i)}-\frac{1}{N}\left(\sum_{k=1}^{N}T_{\rm load}^{(k)}\right)\right|$, which correspond to the differences between the average load torque and the load torques on the original Stirling engines. Depending on the sign of the load torque, the flywheels in system D rotate in a positive or negative direction, but all flywheels perform negative work on the loads (i.e., the load acting on any flywheel applies positive work on the flywheel) regardless of the direction of rotation since the direction of the force exerted by the flywheel pulling the weight is opposite to the direction of the weight’s motion, which is the essential difference from system M. In this sense, the engine ${\rm M}_{\rm EG}$ moves actively w.r.t the load, whereas the flywheels ${\rm D}_{\rm FW}^{(i)}$ are moved passively w.r.t the loads acting on them. When the motions of the flywheels settle to steady states, i.e., when the rotational speeds settle to constant values, the work done by the loads on the flywheels is completely converted to the heat due to viscous friction on the flywheels, which causes the entropy production in system D. If the original Stirling engines are coupled, flywheels ${\rm D}_{\rm FW}^{(i)}$ are also coupled with the same coupling strength and network structure. Unlike the uncoupled case, even if the motions of the flywheels settle to steady states, the rotational speeds are not constant values, so some of the work that the loads and the coupling forces perform to the flywheels are used to accelerate or decelerate the rotating bodies. However, the time-averaged work done by the loads on the flywheels $-\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{D}}\Bigr{\rangle}$ is equal to the time-averaged work that is carried out against the friction torque $\Bigl{\langle}P_{\mathrm{fric}}^{\mathrm{D}}\Bigr{\rangle}$ (cf. Eq.(44) and (45)). Since the coupling between the flywheels attempts to bring the variance of $\omega_{\rm d}^{(i)}$ close to zero, the negative power applied by the flywheels to the loads also approaches zero. Therefore, the averaged brake power of the total system is increased by the weak coupling, and at the same time, the total thermal efficiency is also improved since all the energy from the heat reservoir is absorbed by system M, which is independent of the coupling. When all flywheels in system D stop rotating, which corresponds to the fully synchronized state of the original Stirling engines, both the averaged brake power and the thermal efficiency of the total system reach the maximum values. We thus conclude that the increase of the averaged power and the thermal efficiency of the total system due to synchronous transitions is a result of the fact that the energy dissipation due to the viscous friction in system D is suppressed by synchronous transitions.

### v2

@@ Abstract
This study developed the theory of nonequilibrium thermodynamics for populations of low-temperature-differential (LTD) Stirling engines weakly-coupled in a general class of networks to clarify the effects of synchronous and asynchronous transitions on the power and thermal efficiency. We first show that synchronous (asynchronous) transitions increase (decrease) the power and thermal efficiency of weakly-coupled LTD Stirling engines based on quasilinear response relations between formally defined thermodynamic fluxes and forces. After that, we construct a conceptual model satisfying the quasilinear response relations to give a physical interpretation of the changes in power and thermal efficiency due to synchronous and asynchronous transitions, and justify the use of this conceptual model. We then show that the conceptual model, rather than the quasilinear response relations, preserves the thermodynamic irreversibility of the original model and thus gives more accurate results than those using the quasilinear response relations. Finally, we compare the dynamics between the original and the conceptual models for two-engine systems and show that the conceptual models roughly preserve the dynamical characteristics leading up to the synchronous transitions, while some detailed dynamical structures are lost.
@@ I Introduction
The present study extends the analysis performed in [9] to populations of LTD Stirling engines weakly-coupled in a general class of networks, and constructs the theory of nonequilibrium thermodynamics with synchronous and asynchronous transitions. In Section II, we give a mathematical model of weakly-coupled low-temperature differential Stirling engines. In Section III, we develop the theory of nonequilibrium thermodynamics for the coupled engines as the main results. We first perform the analysis using formally defined thermodynamic fluxes and forces and the quasilinear response relations established between them to show that synchronous (asynchronous) transitions increase (decrease) the power and thermal efficiency. After that, we construct a conceptual model that satisfies the quasilinear response relations to give a physical interpretation of the transitions in power and thermal efficiency, and justify the use of this conceptual model. We then show that the conceptual model, rather than the quasilinear response relations, preserves the thermodynamic irreverbility of the original model and thus gives more accurate results than those using the quasilinear response relations. In Section IV, we compare the dynamics between the original and the conceptual models for two-engine systems and show that the conceptual models roughly preserve the dynamical characteristics leading up to the synchronous transitions, while some detailed dynamical structures are lost. Finally, in Section V, we summarize the study and give the future outlook.
@@ II Model
Here, $\theta$ is the phase of the crank connected to the power piston, and $\omega$ is the angular velocity of that crank. A crank is also attached to the displacer, which is advanced by $\pi/2$ in phase compared to the crank attached to the power piston. These cranks are attached to a flywheel with a large moment of inertia to smoothen the rotational speed of the engine. The physical meanings of the parameters included in the above model are as follows: $\sigma$ is a positive constant determined by the surface areas of the large and small cylinders; $V(\theta)=2+\sigma(1-\cos\theta)$ and $T(\theta,\omega)=T_{\rm eff}(\theta)/\left(1+\frac{\sigma\sin\theta\omega}{GV(\theta)}\right)$ represent the volume and temperature of the gas confined to the cylinders, respectively; $T_{\rm eff}(\theta)=1+\frac{\sin\theta}{2}\Delta T$ is the effective temperature of the heat reservoirs that periodically changes depending on the phase; $G$ is the thermal conductance associated with the heat transfer between the gas and the surface of the large cylinder; $P_{\rm air}$ is the atmospheric pressure acting on the power piston; $\Gamma$ is the friction coefficient associated with the flywheel, and $T_{\rm load}>0$ is the load torque acting on the cranks. The minimal model was obtained by assuming a time-scale separation between the dynamics of the crank and the dynamics of the gas describing the energy conservation law, and the expression for the gas temperature $T(\theta,\omega)$ was obtained by the adiabatic elimination of the fast variable $T$, the gas temperature, in the differential equation describing the gas dynamics [4, 5]. The heat fluxes from the bottom and top surfaces of the large cylinder was assumed to obey the Fourier law
@@
where $G_{\rm m}(\theta)$ with ${\rm m}={\rm b}$ (or t) represents the effective thermal conductance between the gas and the bottom (or top) heat reservoir. It was also assumed that $G_{\rm m}(\theta)\equiv G\chi_{\rm m}(\theta)$, where $\chi_{\rm m}(\theta)\hskip 2.84526pt(0\leq\chi_{\rm m}(\theta)\leq 1)$ is a function that controls the coupling between the gas and the bottom or top heat reservoir, given as $\chi_{\rm b}(\theta)=\frac{1}{2}(1+\sin\theta)$ and $\chi_{\rm t}(\theta)=\frac{1}{2}(1-\sin\theta)$, reflecting the motion of the displacer. The dynamical equations (1a) and (1b) describe the engine as a nonlinear pendulum, where the first term on the right-hand side of Eq. (1b) represents the driven force due to the temperature difference. It has been experimentally verified that the minimal model (1a) and (1b) correctly reflects the essential characteristics of a real LTD Stirling engine [12].
Figure 1: (a) Front view of an LTD Stirling engine. The gas confined to the cylinders is in contact with the bottom and top heat reservoirs with temperatures $T_{\rm b}$ and $T_{\rm t}$. The displacer that advances the power piston in $\pi/2$ phase controls the coupling between the gas and the heat reservoirs. (b) Side view of $N$ LTD Stirling engines coupled in a chain with different load torques acting on the cranks.
@@
Some important assumptions are made about the equations of motion (3a) and (3b) describing the rotational motion and Eq. (4) describing the heat fluxes. We first assume that each engine is in the quasilinear response regime [5] in the absence of coupling, i.e., a limit cycle circling the surface of the phase space $\mathbb{T}\times\mathbb{R}$ that describes the rotational motion exists, and the effective frequency $\langle\omega_{i}\rangle$ and the averaged heat flux $\Bigl{\langle}J_{Q_{\rm b}}^{(i)}\Bigr{\rangle}$ from the heat reservoir of temperature $T_{\rm b}$ show a linear dependency w.r.t. $T_{\rm load}^{(i)}$ and $\Delta T$:
@@
Figure 2: Typical trajectories (red and green plots) of the limit cycle motion of the two engines satisfying condition (9). The parameters of the engines were chosen as follows: $\sigma=0.02$, $p_{\mathrm{air}}=\frac{1}{V(\frac{\pi}{4})}=\frac{1}{2+\sigma(1-\cos\frac{\pi}{4})}\approx 0.49854$, $G=1.5$, $\Gamma=0.001$, $\Delta T=1/29.3$, $T_{\mathrm{load}}^{(1)}=8.5324\times 10^{-7}$, and $T_{\mathrm{load}}^{(2)}=1.2799\times 10^{-5}$.
@@ III Main results: Theory of power and thermal efficiency based on nonequilibrium thermodynamics
In this section, we develop a theory of nonequilibrium thermodynamics for weakly-coupled LTD Stirling engines with synchronous and asynchronous transitions. We first use the quasilinear response relations between thermodynamic fluxes and forces to show that the synchronous (asynchronous) transitions increase (decrease) the power and thermal efficiency in subsection A. After that, we construct a conceptual model satisfying the above quasilinear response relations to give a physical interpretation of the changes in power and thermal efficiency due to the weak coupling in subsection B. We then discuss the justification for using this conceptual model in subsection C. In subsection D, we perform numerical experiments to confirm the validity of the analysis on power and thermal efficiency. Finally, in subsection E, we review the analysis in subsections A -- C and show that the conceptual model, rather than the quasilinear response relations, preserves the thermodynamic irreversibility of the original model and thus gives more accurate results than those using the quasilinear response relations.
@@ III.1 Analysis based on the quasilinear response relations
Here, $\bigl{\langle}P_{\rm m}\bigr{\rangle}\equiv\bigl{\langle}\omega_{\rm m}\bigr{\rangle}\left(\sum_{i=1}^{N}T_{\rm load}^{(i)}\right)$ denotes the power owing to the motion of the mean angle, and $\bigl{\langle}P_{\rm rel}\bigr{\rangle}\equiv\sum_{i=1}^{N}\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}T_{\rm load}^{(i)}$ denotes the power owing to the relative motion. To see the effects of synchronous and asynchronous transitions on the brake power, it is sufficient to investigate the change in $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ since $\bigl{\langle}\omega_{\rm m}\bigr{\rangle}$ is independent of the coupling from the quasilinear response relations. Substituting the quasilinear response relations w.r.t. fluxes $\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$ in Eq. (17), $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ can be rewritten as
@@
We have added a term representing the averaged power due to the weak coupling of the $N$ coupled engines in Eq. (26), which takes the value 00 and is approximated as
@@
and used the quasilinear response relations w.r.t. fluxes $\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$ in Eq. (27). The first equality in equation (28) can be easily derived by noting that $K_{ij}=K_{ji}$; the second approximation comes from Eq. (14) and the third equality can be obtained using Eq. (16). Since $\frac{1}{N}\sum_{i=1}^{N}\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}^{2}$ represents the dispersion of $\langle\omega_{i}\rangle$ around the mean value $\langle\omega_{\rm m}\rangle$, Eq. (27) implies that the coupled system achieves higher averaged brake power with smaller rotational speed dispersion, and that the averaged brake power reaches the maximum when the engines are synchronized. On the other hand, the total heat flux from the heat reservoir of temperature $T_{\rm b}$ is independent of the coupling from the quasilinear response relations (17). This implies that synchronous (asynchronous) transitions improve (deteriorate) both the averaged brake power and the thermal efficiency of the total system.
@@ III.2 Analysis based on the conceptual models constructed from the quasilinear response relations
Equations (31)-(34) indicate that system M can be considered as composed of the same heat reservoirs as those of the original system and a heat engine ${\rm M}_{\rm EG}$ working between the heat reservoirs with viscous damping coefficient of $N\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)$ and load torque $\sum_{i=1}^{N}T_{\rm load}^{(i)}$. Indeed, in this case, by utilizing the energy conservation law
@@
$\biggl{\langle}\frac{d\sigma_{\mathrm{D}}}{dt}\biggr{\rangle}$ $\approx\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)\left(\sum_{i=1}^{N}\biggl{\langle}\left(\omega_{\mathrm{d}}^{(i)}\right)^{2}\biggr{\rangle}\right)$ (43)
@@
Now we use the conceptual model constructed above to illustrate how weak coupling increases the averaged brake power and thermal efficiency of the total system (Note that the conceptual model is refered to as the physical picture of isolated systems M and D, not only the differential equations describing the rotational motion and the heat flux equation. This model was constructed through the quasilinear response relations (17) and thus should reflects the thermodynamic irreversibility of the original coupled Stirling engine system). Figure 3 shows a schematic diagram of systems M and D. The load acting on the engine ${\rm M}_{\rm EG}$ and flywheels ${\rm D}_{\rm FW}^{(i)}$ are represented by weights tied to the rotating bodies. Let us first consider the case where the original Stirling engines are uncoupled. In this case, flywheels ${\rm D}_{\rm FW}^{(i)}$ do not interact either and are simply driven by load torques of magnitude $\left|T_{\rm load}^{(i)}-\frac{1}{N}\left(\sum_{k=1}^{N}T_{\rm load}^{(k)}\right)\right|$, which correspond to the differences between the average load torque and the load torques on the original Stirling engines. Depending on the sign of the load torque, the flywheels in system D rotate in a positive or negative direction, but all flywheels perform negative work on the loads (i.e., the load acting on any flywheel applies positive work on the flywheel) regardless of the direction of rotation since the direction of the force exerted by the flywheel pulling the weight is opposite to the direction of the weight’s motion, which is the essential difference from system M. In this sense, the engine ${\rm M}_{\rm EG}$ moves actively w.r.t the load, whereas the flywheels ${\rm D}_{\rm FW}^{(i)}$ are moved passively w.r.t the loads acting on them. When the motions of the flywheels settle to steady states, i.e., when the rotational speeds settle to constant values, the work done by the loads on the flywheels is completely converted to the heat due to viscous friction on the flywheels, which causes the entropy production in system D. If the original Stirling engines are coupled, flywheels ${\rm D}_{\rm FW}^{(i)}$ are also coupled with the same coupling strength and network structure. Unlike the uncoupled case, even if the motions of the flywheels settle to steady states, the rotational speeds are not constant values, so some of the work that the loads and the coupling forces perform to the flywheels are used to accelerate or decelerate the rotating bodies. However, the time-averaged work done by the loads on the flywheels $-\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{D}}\Bigr{\rangle}$ is equal to the time-averaged work that is carried out against the friction torque $\Bigl{\langle}P_{\mathrm{fric}}^{\mathrm{D}}\Bigr{\rangle}$ (cf. Eq.(43) and (44)). Since the coupling between the flywheels attempts to bring the variance of $\omega_{\rm d}^{(i)}$ close to zero, the negative power applied by the flywheels to the loads also approaches zero. Therefore, the averaged brake power of the total system is increased by the weak coupling, and at the same time, the total thermal efficiency is also improved since all the energy from the heat reservoir is absorbed by system M, which is independent of the coupling. When all flywheels in system D stop rotating, which corresponds to the fully synchronized state of the original Stirling engines, both the averaged brake power and the thermal efficiency of the total system reach the maximum values. We thus conclude that the increase of the averaged power and the thermal efficiency of the total system due to synchronous transitions is a result of the fact that the energy dissipation due to the viscous friction in system D is suppressed by synchronous transitions.